OpenStudy (anonymous):

Which has the same value as (4 + 3i)²

OpenStudy (anonymous):

\[(4 + 3i)²\] \[(4 + 3i)(4 + 3i)\]

OpenStudy (anonymous):

the one that looks like \[(4+3i)(4+3i)=(16-9)+(2\times 4\times 3)i\]

OpenStudy (anonymous):

the possibe answers are A)7 B)7 + 24i C)16 + 9i D)25

OpenStudy (anonymous):

\[(4 + 3i)²\] \[(4 + 3i)(4 + 3i)\] \[16 + 12i + 12i + 9i^2\] \[16 + 24i + 9i^2\] CONTINUE *hint* \[i^2 = -1\]

OpenStudy (anonymous):

@Indiadabest ?

OpenStudy (anonymous):

A) 7 B) 7 + 24i C) 16 + 9i D) 25

OpenStudy (anonymous):

between those answer

OpenStudy (anonymous):

\[\bf (4+3i)(4+3i)=16+24i+9i^2=16+24i+9(\sqrt{-1})^2=..?\] @Indiadabest